OpenStudy (anonymous):

Find the solutions of the equation: F(x) = x^3 – 13x^2 – 42x – 16 Help, please.

OpenStudy (anonymous):

Welcome to OS :) Thats the formula to find F (x), what's the matter?

OpenStudy (anonymous):

I need to find the Zeros or solutions, but i don't know how

OpenStudy (anonymous):

example, x=1 So F(1) = (1)\(^{3}\)-13\(^{2}\)-42-16 =1-169-41-16 =-168-57 = (-225)

OpenStudy (anonymous):

but for it to be right it's supposed to equal 0. do I just go through all possible numbers until then?